### v1

# Vid2Fluid: 3D Dynamic Fluid Assets from Single-View Videos with Generative Gaussian Splatting

Abstract.
The generation of 3D content from single-view images has been extensively studied, but 3D dynamic scene generation with physical consistency from videos remains in its early stages. We propose a novel framework leveraging generative 3D Gaussian Splatting (3DGS) models to extract 3D dynamic fluid objects from single-view videos. The fluid geometry represented by 3DGS is initially generated from single-frame images, then denoised, densified, and aligned across frames. We estimate the fluid surface velocity using optical flow and compute the mainstream of the fluid to refine it. The 3D volumetric velocity field is then derived from the enclosed surface. The velocity field is then converted into a divergence-free, grid-based representation, enabling the optimization of simulation parameters through its differentiability across frames. This process results in simulation-ready fluid assets with physical dynamics closely matching those observed in the source video. Our approach is applicable to various fluid types, including gas, liquid, and viscous fluids, and allows users to edit the output geometry or extend movement durations seamlessly. Our automatic method for creating 3D dynamic fluid assets from single-view videos, easily obtainable from the internet, shows great potential for generating large-scale 3D fluid assets at a low cost.
Figure 1. Vid2Fluid takes a single-view video of fluid as input shown on the left and automatically extract high-quality ready-to-use dynamic fluid asset shown on the right. In case (a), we preserve the camera viewpoint and the appearance of the original fluid while in case (b) the appearance and the viewpoint are modified.
1. Introduction
The field of 3D generation has gained significant attention in recent years, with researchers exploring innovative ways to integrate traditional 3D modeling applications with cutting-edge 3D generation technologies. It aims to enhance design and editing inspiration while reducing labor costs, paving the way for more efficient and creative digital assets production. By combining 3D generation with various editing, animation, and simulation techniques, a new paradigm for digital content creation is emerging. One of the most challenging problems is combining 3D generation with physics-based simulation, particularly in the context of fluid dynamics. While most existing research focuses on rigid body generation and simulation due to its more stable and predictable outcomes, as well as the relative simplicity of rigid body motion compared to the complex behaviors exhibited by fluids , the demand for generative physics asset creation for fluids remains high. Because the manual creation of fluid assets is not only time-consuming but also labor-intensive, primarily due to the intricate nature of fluid dynamics and the need for precise control over various simulation parameters.
Despite the growing demand, the research on fluid 3D generation and simulation remains relatively unexplored. To address this need, we have conducted extensive research to understand the requirements and pain points of artists and content creators in the field. Our findings reveal several key insights: (1) Artists desire a tool that can directly generate fluid assets from a given video, preserving the visual and dynamic characteristics of the original content. (2) In most cases, artists are interested in fluid assets with relatively small advancement, which simplifies the problem and reduces the complexity of the simulation. (3) The generated fluid assets should be easily editable and integrable into existing 3D scenes.
Based on these insights, we proposed a novel method, Vid2Fluid, that addresses the unique challenges of extracting 3D dynamic fluid assets from single-view videos. We approximate the fluid with a meaningful geometry represented by 3D Gaussians (3DGS) based on generative methods and ensure the geometry consistency across the frames. The volumetric field is then estimated based on the 3D generated geometry and motion presented in the video. We also design an optimization framework to approximate the simulation parameters of the fluid. Our experiments show that our method produces high-quality results for fluids with normal advancement, making it suitable for a wide range of artistic and practical applications. Moreover, our method also demonstrate high versatility through extensive experiments, showcasing its effectiveness in handling various types of fluids, rendering options, and editable features.
Our main contributions can be summarized as follows:
  • The problem we aim to solve is a novel task with practical real-world applications and user-editable features. We will open-source the implementation of our method.
  • We design the Geometry and Motion Reconstruction module to unify the inconsistent 3DGS generated fluid from different frames of the video into a coherent geometry and accurately estimate the volumetric velocity field of the fluid.
  • We realize a differentiable grid velocity evolution procedure to optimize the simulation parameters that are compatible with general grid-based or hybrid simulation methods.
Figure 2. Overview of our framework.
2. Related Work
2.1. 3D Generation with Gaussian Splatting
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) adopts a point-based radiance field, using 3D Gaussian primitives to represent scenes. It has emerged as a prevalent research topic in 3D representation (Zhang et al., 2024b; Huang et al., 2024; Yu et al., 2024b; Lu et al., 2024; Yu et al., 2024a) due to its ability to depict high-quality geometry and textures in novel view synthesis.
3DGS provides a new perspective not only for real-time scene reconstruction but also for 3D generation. DreamGaussian (Tang et al., 2023) optimizes a 3DGS through score distillation using a pre-trained text-to-image diffusion model. Recent methods (He et al., 2025; Zhang et al., 2024a; Zhou et al., 2024) investigate directly train diffusion models on Gaussian splats for higher efficiency. However, the direct methods may struggle to handle real-world inputs, like fluid objects, since they are usually trained on synthetic 3D datasets. LGM (Tang et al., 2025) transforms the single-view generation problem into a multi-view 3DGS reconstruction task using a pre-trained single-view to multi-view 2D diffusion model. The most recent transformer-based methods (Zou et al., 2024; Xiang et al., 2024) are proposed to achieve faster generation with higher quality. TriplaneGaussian (Zou et al., 2024) creates a point cloud from a single image and uses a hybrid triplane-Gaussian representation to greatly accelerate the generation. TRELLIS (Xiang et al., 2024) employs rectified flow transformers unified structured latent to get high-quality results.
2.2. Physics-based Fluid Simulation and Dynamic Gaussians
In physics-based simulation, fluids are often simulated using Eulerian grids and/or Lagrangian particles (Bridson, 2015). Resolving fluid dynamics requires solving the simplified version of the Euler equation (Bardos and Titi, 2007) or the complete version of the Navier-Stokes equation (Doering and Gibbon, 1995) regarding viscosity. Methods of solving these equations advanced from giving analytic solutions  (Wejchert and Haumann, 1991; Stam and Fiume, 1993) to later numerical discretization schemes  (Foster and Metaxas, 1996; Tome and McKee, 1994). On the Lagrangian side, Smoothed Particle Hydrodynamics (SPH) methods  (Monaghan, 1992; Becker and Teschner, 2007; Solenthaler and Pajarola, 2009; Ihmsen et al., 2014; Bender and Koschier, 2015) could simulate amorphous objects and particle systems for sub-grid movements. Alternatively to solving force-based equations, Position-based Dynamics  (Müller et al., 2007; Macklin et al., 2016; Macklin and Müller, 2013) was introduced to constrain object positions directly and achieve higher simulation efficiency. For simulation stability, Stam  (Stam, 1999) introduces the concept of semi-Lagrangian advection which brings up the idea of the hybrid field. Hybrid methods  (Harlow, 1988; Brackbill and Ruppel, 1986; Jiang et al., 2015; Fang et al., 2020; Jiang et al., 2016; Hu et al., 2018) combine advantages of Lagrangian and Eulerian schemes, representing fluid by particles while computing dynamics on grids.
The neural radiance expression of objects with dynamics has long been studied for NeRF systems, including deformation capture by canonical and displacement fields  (Park et al., 2021a, b) and dynamics 3D synthesis  (Gao et al., 2021; Li et al., 2021a; Liu et al., 2022; Pumarola et al., 2021; Qiao et al., 2023). Similarly, for the later proposed 3D Gaussian Splatting, dynamics properties have been imposed on the explicit radiance representation of Gaussians. PhysGaussian (Xie et al., 2023) integrated MPM framework to enable reconstructed 3D Gaussians with versatile dynamic behaviors, and GaussianSplashing (Feng et al., 2024) combine Position-based Dynamics with advanced rendering methods to represent Gaussian particles with fluid dynamics and appearances. Instead of giving physical parameters to reconstructed Gaussian subjectively, several works have developed the learning procedure of 3DGS dynamics from references. PhysDreamer (Zhang et al., 2024c) optimizes material parameters through a differentiable MPM simulation to reproduce dynamic Gaussians that behave similarly to input videos. BAGS (Peng et al., 2024) learns the weights of imposed Gaussian Ellipsoid Neural Bounds to animate the reconstructed object according to the input videos. In our work, we embed generated Guardians with physical parameters compatible with APIC method, to align the dynamics behavior of simulated Gaussian particles closest to that of the fluid in the video.
2.3. Velocity Extraction from Videos
We limit the discussion to fluid velocity reconstruction from input videos and images, instead of direct velocity data or experiments. To collect pixel information from videos, optical flow could be effectively utilized in neural reconstructions  (Li et al., 2021b; Du et al., 2021) for general flowing scenes that vary for image styles and object appearances. Targeting on specific fluid textures can noticeably improve learning of dynamics and re-simulation accuracy  (Guan et al., 2022; Liu et al., 2023). For single-view image sequences, vision techniques like tomography can be utilized with some physical priors to achieve visual processing that retains geometric and physical consistency with input fluids (Gregson et al., 2014; Zang et al., 2020). Generally, multi-view inputs are required for 3D velocity reconstruction and novel view fluid synthesis relies on at least sparse viewpoints of the fluid for trained neural networks to infer the unseen sides  (Okabe et al., 2015; Eckert et al., 2018, 2019). Thus, previous work seldom focuses on reconstructing the entire velocity field from the single input as a generative process. To inform the extracted velocity field with physical grounded features, DVP (Deng et al., 2023) successfully encodes vortex features to learn the specific eddy dynamics of fluid but is restricted to 2D profile. Other works  (Franz et al., 2021; Chu et al., 2022; Yu et al., 2024c) extend to 3D space with strict physical constraints and volume rendering, achieving high-fidelity reconstruction while limited to certain fluid types. In our work, we use single-view images to reconstruct the 3D velocity field even for novel views, leveraging the idea of image-conditioned 3D generation.
3. Method
3.1. Overview
Vid2Fluid takes a single-view video of fluid as input and outputs a 3D fluid physics asset. We design a two-stage pipeline to automatically extract crucial information, like the fluid’s geometry, appearances, physical properties, and motion, from the input video to form the final fluid physics. Stage 1: Geometry and Motion Reconstruction. We employ image-conditioned 3D generative models to generate 3DGS for the input frames and preprocess the generated 3DGSs to make them unified, denoised, and dense in the volume. With the 3D geometry information, we estimate the volumetric velocity field of the fluid by interpolating the surface velocity extracted from the frame images into the volume (Sec. 3.3). Stage 2: Simulation Parameter Optimization. To estimate the fluid dynamics in the video through physical evolution, we convert the velocity field on points to a grid-based representation to enable the differentiable grid velocity computation and guide the optimization of simulation parameters, including fluid physical properties and boundary conditions (Sec. 3.4).
3.2. Background
3.2.1. Generative 3D Gaussian Splatting
3D Gaussian Splatting is an explicit radiance-based representation of 3D objects, utilizing high-degree features of shape and color for multi-view synthesis. Due to its differentiable volume rendering capabilities, it can be effectively employed in image-to-3D object generation, enabling the alignment of conditioned image textures. This generation procedure, denoted as $G$, can be described as: $G:\{C\}_{w,h}\rightarrow{\{x,\sigma,A,F\}}_{p}$, where $\{C\}_{w,h}$ represents pixels of the conditioned image at position $(w,h)$. $x$, $\sigma$, $A$, and $F$ denote the position, opacity, covariance matrix, and spherical harmonic features of each Gaussian particle $p$, respectively. Ideally, the input image could be recovered through discretized splatting rendering:
(1) $\sum_{i\in\mathcal{N}}\alpha_{i}\boldsymbol{SH}(r|_{w,h};F_{i})\displaystyle\prod_{j=1}^{i-1}(1-\alpha_{j})\rightarrow\{C\}_{w,h},$
where $\alpha_{i}$ is the product of $\sigma_{i}$ and the projected 2D Gaussian density of where the kernel intersects with the ray in direction $r|_{w,h}$ from the specific pixel at $(w,h)$, and $\boldsymbol{SH}$ denotes the color calculated with features $F_{i}$. This is equivalent to computing a depth-and-opacity weighted average color of particles along the ray direction. Consequently, individual Gaussian points possess adequate geometric and chromatic information, facilitating their representation of point-based models, such as fluids.
3.2.2. Affine Particle in Cell Method and Dynamic 3D Gaussians
Affine Particle-in-Cell (APIC) (Jiang et al., 2015) is a numerical technique for simulating particles in continuous media, enhancing stability and precision through additional computation of local affine velocity fields. In a time step $n$, the particle velocity is first transferred to the grid together with its affine part:
(2) $(mv)_{i}^{n}$ $=\sum_{p}w_{ip}m_{p}[v^{n}_{p}+\mathbf{C}_{p}^{n}(\mathbf{x}_{i}-\mathbf{x}_{p}^{n})]$
$v_{i}^{n}$ $=(mv)_{i}^{n}\big{/}\sum_{p}w_{ip}m_{p},$
where $i$ and $p$ denote grid and particle coordinates respectively. $m$ is the mass, $v$ is the velocity, $\omega$ is the weight function, such as quadratic B-Spline kernel, $\mathbf{C}$ is the affine matrix associated with each particle. Assuming the simulated media is incompressible, the velocity change $\frac{\partial v}{\partial t}$ can be computed on the grid by the following equation:
(3) $\frac{\partial v}{\partial t}=$ $-\frac{\nabla p}{\rho}+\nabla\cdot(2\nu\textbf{S}_{ij})+g$
$\textbf{S}_{ij}=$ $\frac{1}{2}[\nabla\emph{v}+(\nabla\emph{v})^{\mathsf{T}}],$
[…]
(4) $\mathbf{v}_{p}^{n+1}$ $=\sum_{i}w_{ip}(\mathbf{v}_{i}^{n}+\frac{\partial v^{n}_{i}}{\partial t}),$
$\mathbf{C}_{p}^{n+1}$ $=\frac{4}{\Delta x^{2}}\sum_{i}w_{ip}\mathbf{v}_{i}^{n+1}(\mathbf{x}_{i}-\mathbf{x}_{p}^{n})^{T}~{}(\text{quadratic kernel}).$
Finally, the convective effect of simulated media is achieved on the particles by $\mathbf{x}_{p}^{n+1}=\mathbf{x}_{p}^{n}+\mathbf{v}_{p}^{n+1}\Delta t$.
3.3. Reconstruct 3D Gaussians with Dynamics
3.3.1. Preprocess with Generative Gaussian Splatting
Since our input only contains single-view information, to obtain the 3D
information, we rely on the existing single-image to 3DGS generation methods (Zou et al., 2024; Xiang et al., 2024). However, these methods are trained on datasets of 3D surface representation and the generated 3D Gaussians are typically sparse and concentrated on the object’s surfaces. The generated 3DGS need to be denoised and densified so that the Gaussian points could precisely carry every pixel information from the frame. As shown in the inset, we sample the 3DGS space into grids and insert a Gaussian at the center of the grid $p_{c}$ if $p_{c}$ is inside the fluid by checking if over half of the number of the intersections of 3DGS’s outer hull and the rays from $p_{c}$ is odd. Additionally, the Gaussian primitives with low opacity or heavily stretchered covariance are pruned.
We also need to carefully deal with the consistency of geometry for the generated 3D Gaussians in continuous frames. Generative methods do not guarantee that the output variation is on the same scale with the perturbation of input conditions. Moreover, adjacent frames could be generated with quite different geometries at the backside from the camera, though the foreground is conditioned by similar inputs. This is acceptable for free-moving fluid like smoke, which shows highly dynamic behavior in the video. However, this inconsistency can impact the calculation of physically grounded fluid dynamics for flowing rivers on fixed riverbeds. To address the issue, we perform the union operation on the generated 3D Gaussians from $N$ consecutive frames, forming a batch to be used in the subsequent process of simulation parameter optimization in Sec. 3.4. Note that $N$, the number of frames to union, is dynamically decided by the moving intensity of the object profile. This is achieved by evaluating similarities between adjacent frames $f$ and $f+1$ throughout the video: $N\propto MSE\big{(}PSNR({f,f+1})\big{)}$.
3.3.2. 3D Fluid Surface Velocity Estimation
After obtaining the geometry, we can estimate the volumetric velocity field of the object. This starts by first estimating the 3D surface velocity, where we utilize optical flow (Farnebäck, 2003) in the screen space, as pixels moving on the screen can be seen as material points being displaced in the Normalized Device Coordinates (NDC). Additionally, we combine the $z$-depth information from the preprocessed 3DGS to get displacements in the third dimension. However, optical flow heavily relies on the color change of pixels between frames, which is not detectable in the area that exhibits a uniform flow, where no splashing or vortex effects could be exploited. To tackle this problem, we propose a mainstream direction detection method and compensate the region where optical flow fails. The mainstream direction is determined by edge orientations of fluid detected from the video, and then mapped to the entire region of fluid. When filling in the missing parts of the optical flow result, we project neighboring optical flows in the mainstream direction, with those more aligned with the main direction receiving greater weighting.
(5) $\vec{d}_{opz,k}=\sum_{i\in B(k)}w(i)\max(0,\bigg{(}\vec{n}_{k}\cdot\frac{\vec{d}_{opz,i}}{|d_{opz,i}|}\bigg{)})\cdot\vec{d}_{opz,i},$
where $\vec{d}_{opz}$ is the 3D velocity in NDC space, calculated from optical flow and depth change. $i$ denotes the detectable place by optical flow, and $k$ is the missing part. $B$ bounds the range of $i$ contributing to specific $k$ and defines certain distance-based weight $\omega(i)$. The dot product of mainstream direction $\vec{n}_{k}$ and the normalized $\vec{d}_{i}$ measures the cosine value of their angle.
We then project the velocity on the screen space to the 3D Gaussian space: $\vec{v_{p}}=\big{(}S\circ(PW)_{+}T\big{)}^{-1}\circ\vec{d}_{opz,i}$ , where $\vec{v}_{p}$ is the particle velocity, $P$ is the view projection matrix, and $W$ is the world-to-camera matrix. $T$ and $S$ denote the transition and scaling same as those (if any) on input images before generating the 3DGS.
$3$D Volumetric Velocity Estimation
Figure 3. Illustration (left) and a real volumetric velocity field (right). The velocity is dampen near the riverbed. On the right figure, the darker the color, the smaller the velocity.
With the surface velocity extracted, we interpolate it across the entire volume of preprocessed Gaussian particles to initialize the dynamic behavior of the fluid object. Before interpolation, the enclosure of the fluid must be determined, especially in the regions of 3D Gaussians that are unseen from the camera. Depending on the constraint strength of the boundary, we let the surface velocity determine relative magnitudes and impose the presumed velocity profile. For flowing liquid, the boundary layer approximation is imposed on the riverbed as shown in Fig. 3, but for freely moving fluid, we do not dampen the velocity on the unseen sides. This ensures a gradual change in the velocity of the flowing layer as it approaches any solid boundary, where the velocity should ultimately be zero. To reveal this effect, we choose the third-order polynomial of the integral approximate solution of laminar external flows:
(6) $\frac{|\vec{v}|}{V_{\infty}}=\frac{3}{2}\left(\frac{y}{\delta}\right)-\frac{1}{2}\left(\frac{y}{\delta}\right)^{3},$
where $V_{\infty}$ is the magnitude of corresponding surface velocity, $\delta$ is the boundary layer thickness, which is dynamically decided by the fluid type: for a flowing river it is set to a larger thickness, while for smoke, it is limited to a sub-cell scale so it introduces very small damping. $\vec{v}$ takes the direction of the mainstream but not the same as its surface counterpart. After enclosing the surface of the object, we calculate the inner velocity field by interpolation weighted as a Gaussian distance distribution and finalize it by another divergence-free projection in the next step.
3.4. Presumed APIC Simulation and Optimization
3.4.1. Grid-Based Velocity Evolution
We transfer the velocity field from particles to the grid, which offers two main advantages. First, grid-based representation helps us more efficiently track the fluid movement between two frames. Second, by solving Eq. 3 alongside grid-based convection, particle velocity can be transferred to grid using Eq. 2 with $\mathbf{C}$ set to zero, making the dynamics of Gaussians differentiable.
To formulate the optimization problem , we modify two parts of Eq. 3. First, to solve the Poisson Pressure Equation, instead of using iterative methods such as Jacobi iterations, we implement and modify the PeRCNN (Rao et al., 2021) structure shown in the inset where $P^{(k)}$, $B.C.$ and $\vec{v}$ are the latents of pressure , boundary condition and velocity respectively at $k$-th recurrent layer.
The approach is as accurate as the numerical methods while facilitates the convergence of the optimization. Secondly, as the original APIC simulation method does not solve the convection part on the grid, we additionally integrate the back-tracing mechanics from stable flow (Stam, 1999) in the forward process, allowing for unconditional stability in resolving grid-based velocity. However, since back-tracing involves grid indexing and round-off operations, which reduce the differentiability of the optimizing process, we directly use the analytic gradient of convection term of the split PDE: $\frac{\partial\vec{v}}{\partial t}=-\vec{v}\nabla\vec{v},~{}\frac{\partial\vec{v}}{\partial v_{i}^{{}^{\prime}}}=-dt\frac{\partial\vec{v}}{\partial i},$ where $\vec{v}$ denotes the grid velocity after convection, $v_{i}^{{}^{\prime}}$ denotes the velocity component of dimension $i$ before convection.
3.4.2. Loss Design and Parameter Optimization
Prepared with the differentiable computation of grid velocity, we can optimize several parameters to provide preprocessed 3D Gaussians with physical properties, including bulk velocity at the inlet $v_{in}$ and outlet $v_{out}$, velocity fluctuations at the inlet $\tilde{v}_{in}$, density $\rho$ and viscosity $\nu$ of the fluid, bouncing $b$ and damping $d$ coefficients at the boundary, gravity $g$, and the time step $dt$ of the simulation. To design a reasonable loss function for the optimization, we combine the $L_{2}$ loss of grid velocities and the dot product of normalized velocities as the final loss, masked by grid types of surface, occupation, or empty.
(7) $\vec{v}_{sim}^{i+n}=\text{DiffSim}^{n}$ $\big{(}\vec{v}^{i}_{gt};v_{in},\tilde{v}_{in},v_{out},\rho,\nu,(b,d)|_{{}_{B.C.}},g,dt\big{)}$
$\min~{}\text{gridMask}\circ$ $\bigg{(}A(\frac{\vec{v}_{sim}}{|v_{sim}|}\cdot\frac{\vec{v}_{gt}}{|v_{gt}|})+BL_{2}(\vec{v}_{sim},\vec{v}_{gt})\bigg{)}.$
3.4.3. Augmentation of Optimized Result
In the post-processing of the generated 3D Gaussians and optimized simulating parameters, we can re-simulate the Gaussians using the APIC method, combining Lagrangian and Eulerian spaces. When integrating the APIC with Gaussian points, the simulation properties and Gaussian properties are in two-way coupling. On one hand, Gaussian points determine the initial position and possibly the velocity of the simulating particles. On the other hand, the calculated affine matrix records the local deformation of the material and updates the Gaussian covariance matrix:
(8) $F_{p}^{n+1}$ $=(I+dtC_{p}^{n+1})F_{p}^{n},$
$A_{p}^{n}$ $=F_{p}^{n}A_{p}^{0}F_{p}^{T}.$
Here we approximate the affine matrix $C$ as the velocity gradient and update the deformation gradient $F$ each time step, which then deforms 3D Gaussians and is reflected by the change of its covariance $A$.
4. Results
We implement the framework of Vid2Fluid mainly using Python on a workstation PC equipped with a 24-core Intel(R) Xeon(R) Platinum 8255C CPU and an NVIDIA Tesla V100 GPU. For the 3DGS generation part, we port the open-source implementation of the existing methods (we use TriplaneGaussian (Zou et al., 2024) by default), and output intermediate states. For re-simulating the flowing fluid, we implemented a GPU-parallelized APIC method with Taichi (Hu et al., 2019). The rendering of obtained 3D Gaussians is compatible with a general 3DGS visualizer. We also implement a relighting interface using GaussianShader (Jiang et al., 2023) to achieve higher resolution and incorporate transparent fluid rendering within the original 3DGS CUDA rasterizer.
4.1. Ablation Studies
4.1.1. Denoising and Filling of 3D Gaussians
For the image-conditioned 3DGS, downsampling (e.g., reducing the image resolution to $256\times 256$ pixels in TriplaneGaussian) results in sparse Gaussian points, each with a large elliptical shape. This is undesirable when relating continuous surface velocities with such sparse 3D Gaussians. Thus, as shown in Fig.4, we preprocess the generated Gaussians, which leads to a more compact and accurate representation as mentioned in Sec. 3.3.
Figure 4. Proprecessing operations. The denoising operation removes outlier Gaussian points, returning a smoother geometry (left bottom) compared the geometry on left top. The filling operation inserts Gaussians into sparse generated 3DGS (middle top) and output a dense 3DGS (middle bottom). Our union strategy merges the generated 3DGS from multiple frames and outputs a higher-fidelity geometry (right bottom) compared direct union (right top).
4.1.2. Mainstream-corrected Optical Flow
As mentioned in Sec. 3.3.2, optical flow may return a 2D velocity field with “vacant” static regions. Fig. 5 shows that the mainstream-corrected result accurately captures the true motion where the original optical flow detects near-zero velocities.
Figure 5. The mainstream guided optical flow result (right) presents a more meaningful velocity field compared the unguided result where the velocity is vanished in regions such as the one in the closeup (left).
4.1.3. Neural Surrogate for Poisson Pressure Equation
Fig. 6 demonstrate that our modified PeRCNN (Rao et al., 2021) architecture for solving the Poisson pressure equation is equivalently accurate compared with traditional numerical iterative method. Also, the comparison between the optimization processes of PeRCNN and Jacobi iterations reveals that PeRCNN achieves faster convergence.
Figure 6. Left: We demonstrate that our PeRCNN-based optimization converges in the case of honey. Right: We compared our method with the traditional numerical method to show that our method achieves faster convergence.
4.2. Evaluation on Different Fluid Types
We apply our method to three different types of fluids – gas, inviscid liquid, and viscous fluid – to show the generalization of our framework. For the gas case, we compared our method with HyFluid (Yu et al., 2023) in terms of reconstruction quality and novel view synthesis. It is important to note that HyFluid requires multi-view images as input, whereas our approach relies on only single-view images. As shown in Fig. 7, our result is visually more similar to the input image, thanks to the high-degree harmonic features of 3DGS that align effectively with the input image. In terms of novel views, our framework incorporates a generative process, while HyFluid directly takes as input images from different viewpoints, which allows it to achieve higher similarity to the true smoke in these cases.
Figure 7. Comparison with Hyfluid. Top: Input video frames. Middle: Hyfluid output. Botton: Our output dynamic fluid.
The effects of gravity and boundary conditions are more pronounced in fluids with well-defined shapes. For viscous fluids, the relative motion between different parts is constrained by viscosity and further dampened by the boundary. As a result, by using the same governing equations with different scales of optimized simulation parameters, the reconstructed fluid can effectively mimic a wide range of dynamics observed in the video as shown in Fig. 10.
4.3. Video Velocity Reconstruction
Two synthetic videos from known fluid data are used to compare the reconstructed 2D velocity fields with ground truth. As Vid2Fluid inputs planar velocity that is discretized by frame rate and scaled to screen dimensions, both ground truth and reconstructed 3D velocities are projected into 2D NDC space for comparison for visual examination. Figure 8 shows results for two fluid types. The reconstructed in-plane velocity can capture meso-scale profiles but retains noises on fluid surfaces. Optimizing physics-based simulation and aligning velocity direction yields smoother results and enhances physical details, such as boundary collisions in the river and strain development in honey inflow.
Figure 8. Comparison of planar velocity for fluid as river (top) and honey (bottom). Left column is the ground truth, middle column is the result by Vid2Fluid, and right column is the estimated planar velocity in early stage of the framework as described Fig.2.
4.4. Performance with Different 3DGS Generation Methods
As a generic framework based on generative 3DGS, Vid2Fluid is not dependent on a specific method of Gaussian generation. We test our method with different 3DGS generation methods: generative methods, TriplaneGaussion (TriGS) and TRELLIS (Xiang et al., 2024), and multi-view reconstruction method, vanilla 3DGS (Kerbl et al., 2023). These methods differ in terms of image fidelity, resolution, and creativity, thus catering to varying user needs in real-world applications. As shown in Fig. 11, multi-view optimized 3DGS exhibited the highest quality and fidelity with the input images, but its usage in real-world scenarios is limited. TriGS benefits from triplane encoding of image textures, which aligns well with the input image, though this comes at the cost of lower resolution. TRELLIS shows more divergence from the input images, even at high Classifier Free Guidance (CFG) scales, but produces Gaussians at a higher resolution compared to TriGS.
4.5. Different Rendering Options
The output 3D Gaussians dynamic assets can be rendered mainly in two ways: retaining the originally generated spherical harmonic features (like the mountain river example rendered in Fig. 11 bottom) or reconstructing meshes for high-resolution rendering and interaction (the mountain river example in Fig. 1 top). Without assigning uniform appearances to fluid, 3D Gaussian rasterization relies on the unknown spherical harmonic features of inlet particles. Thus, a natural choice is to sample features from particles at the initial position. We note that the generative method often produces Gaussians at low resolution, which may limit the application of harmonic features, depending on the fluid type.
Unlike a closed system, our fluid assets include inlet particles, which complicates the prediction of inflow textures. While the appearance of the fluid typically aligns with uniform flow, it may lack consistency with the reconstructed appearance. To enable interaction between the generated Gaussians and other assets, users might need to re-render the fluid using the same illumination pipeline as the rest of the scene. For compatibility with splatting rendering, we could adapt the existing 3D Gaussian relighting frameworks (Jiang et al., 2024) and implement the Gaussian fluid with transparent light absorption  (Feng et al., 2024), ensuring rendering efficiency with 3DGS. Moreover, with current GPU-accelerated parallelization, the simulation process achieves real-time speeds, enabling both rasterization and simulation of Gaussians on the fly.
4.6. User Editing
For users looking to edit the generated fluid asset, the framework provides easy modification of simulation parameters to achieve different fluid effects, change the fluid’s material, adjust the downstream flow, or introduce solid-liquid interactions as shown in Fig. 12. In this way, fluid behavior within scenic videos can be faithfully reproduced in the digital world.
4.7. Runtime
We analyze the running time consumed by our method. The results indicate that in most cases, simulation parameter optimization dominates the main runtime as the proportion shown in Fig. 9. We note that in the optimization stage, the runtime is influenced by two key factors: the length of the input video and the grid density.
Figure 9. Runtime (in second) of the example shown in Fig. 1 (b).
5. Conclusions
In this paper, we introduced a novel two-stage pipeline for reconstructing consistent and accurate fluid dynamics from a single-view video, starting with geometry and motion reconstruction and then optimizing simulation parameters. This approach has opened up new possibilities in the field of fluid asset creation. However, our approach also has certain limitations. Firstly, it currently cannot handle long videos with high-torrent fluids due to increased complexity and convergence issues in simulation parameter optimization. Color changes not related to velocity (e.g., in lava due to temperature) may also affect the accuracy of our optical-flow-based velocity estimation. In addition, we note that an optional requirement for our method is that camera parameters are needed for single-view videos with a moving camera. Future work will aim to address these limitations for more accurate reconstructions.