### v2

# 3D Dynamic Fluid Assets from Single-View Videos with Generative Gaussian Splatting

Abstract
While the generation of 3D content from single-view images has been extensively studied, the creation of physically consistent 3D dynamic scenes from videos remains in its early stages.
We propose a novel framework leveraging generative 3D Gaussian Splatting (3DGS) models to extract and re-simulate 3D dynamic fluid objects from single-view videos using simulation methods.
The fluid geometry represented by 3DGS is initially generated and optimized from single-view images, then denoised, densified, and aligned across frames.
We estimate the fluid surface velocity using optical flow, propose a mainstream extraction algorithm to refine it.
The 3D volumetric velocity field is then derived from the velocity of the fluid’s enclosed surface.
The velocity field is therewith converted into a divergence-free, grid-based representation, enabling the optimization of simulation parameters through its differentiability across frames. This process outputs simulation-ready fluid assets with physical dynamics closely matching those observed in the source video.
Our approach is applicable to various liquid fluids, including inviscid and viscous types, and allows users to edit the output geometry or extend movement durations seamlessly. This automatic method for creating 3D dynamic fluid assets from single-view videos, easily obtainable from the internet, shows great potential for generating large-scale 3D fluid assets at a low cost.
[…]
volume: 45issue: 2
1 Introduction
The field of 3D generation has gained significant attention in recent years, with researchers exploring innovative ways to integrate traditional 3D modeling applications with cutting-edge 3D generation technologies. It aims to enhance design and editing inspiration while reducing labor costs, paving the way for more efficient and creative digital assets production. By combining 3D generation with various editing, animation, and simulation techniques, a new paradigm for digital content creation is emerging. One of the most challenging problems is combining 3D generation with physics-based simulation, particularly in the context of fluid dynamics. While most existing research focuses on rigid body generation and simulation due to its more stable and predictable outcomes, as well as the relative simplicity of rigid body motion compared to the complex behaviors exhibited by fluids , the demand for generative physics asset creation for fluids remains high. Because the manual creation of fluid assets is not only time-consuming but also labor-intensive, primarily due to the intricate nature of fluid dynamics and the need for precise control over various simulation parameters.
Despite the growing demand, the research on fluid 3D generation and simulation remains relatively unexplored. To address this need, we have conducted extensive research to understand the requirements and pain points of artists and content creators in the field. Our findings reveal several key insights: (1) Artists seek a tool that can directly generate fluid assets from a given video, preserving the original geometry and dynamic characteristics while enabling customizable digital rendering. (2) In most cases, artists are interested in fluid assets with relatively small advancement, which simplifies the problem and reduces the complexity of the simulation. (3) The generated fluid assets should be easily editable and integrable into existing 3D scenes.
Based on these insights, we proposed a novel method that addresses the unique challenges of extracting 3D dynamic fluid assets from single-view videos through physics-based simulation. We approximate the fluid with a meaningful geometry represented by 3D Gaussians (3DGS) based on generative methods and ensure the geometry consistency across the frames. The volumetric field is then estimated based on the 3D generated geometry and motion presented in the video. This process leverages optic flow to analyze the pixel motion in the video, corrected by a physics-derived constraint to retain real fluid features. We also design an optimization framework to approximate the simulation parameters of the fluid that best recovers the dynamics from the video. Our experiments show that the proposed method produces high-quality results for fluids with convectional flow motion, making it suitable for a wide range of artistic and practical applications. Comparisons are demonstrated to show the necessity and efficacy of the optimization process, whereas simply reconstructing precise geometry but manually simulating the fluid produces far inferior results. Moreover, our method also demonstrate high versatility through extensive experiments, showcasing its effectiveness in handling various types of fluids, rendering options, and editable features.
It should be noted that our method does not prioritize precise reconstruction of the visual appearance. First, the target scene is an open system with inflow and outflow, which is different from closed systems like deformable solids where per-vertex tracking is feasible. Second, our single-view input relies on generative inpainting to complete the 3D appearance, where the limited-view guidance provides no information about other perspectives. Finally, the amorphous nature of fluids means the exposure of the inner part easily breaks the surface texture which is more reliably inferred from surface physics. Therefore, we focus more on recovering fluid dynamics by re-simulation to reveal its underlying physics. Implicit representations of fluid motion often fail to preserve physical laws such as incompressibility and momentum conservation. These high-degree-of-freedom phenomena are poorly represented by directly optimizing displacement or velocity fields, which also lack temporal scalability. In contrast, our physics-based approach can generate spatially and temporally coherent 3D fluids that closely match the video dynamics and are also extendable. Moreover, it provides digital creators with an editable and interactive model, as the intuitive physical parameters are straightforward to tune.
The main contributions can be summarized as follows:
  • We propose a novel open-source framework to extract 3D fluid assets from single-view videos, tackling an under-explored problem with practical real-world applications and user-editable features.
  • We design customized geometry and motion reconstruction strategies to produce more coherent geometry and physically faithful velocity field.
  • We realize a differentiable grid velocity evolution procedure to optimize the simulation parameters that are compatible with general grid-based or hybrid simulation methods.
2 Related Work
2.1 3D Generation with Gaussian Splatting
3D Gaussian Splatting (3DGS) [KKLD23] adopts a point-based radiance field, using 3D Gaussian primitives to represent scenes. It has emerged as a prevalent research topic in 3D representation [ZFS24, HYC24, YSG24, LYX24, YCH24] due to its ability to depict high-quality geometry and textures in novel view synthesis.
3DGS provides a new perspective not only for real-time scene reconstruction but also for 3D generation. DreamGaussian [TRZ23] optimizes a 3DGS through score distillation using a pre-trained text-to-image diffusion model. Recent methods [HCP25, ZCY24, ZZL24] investigate directly training diffusion models on Gaussian splats for higher efficiency. However, the direct methods may struggle to handle real-world inputs, like fluid objects, since they are usually trained on synthetic 3D datasets. LGM [TCC25] transforms the single-view generation problem into a multi-view 3DGS reconstruction task using a pre-trained single-view to multi-view 2D diffusion model. The most recent transformer-based methods [ZYG24, XLX24] are proposed to achieve faster generation with higher quality. TriplaneGaussian [ZYG24] creates a point cloud from a single image and uses a hybrid triplane-Gaussian representation to greatly accelerate the generation. TRELLIS [XLX24] employs rectified flow transformers unifying structured latent to get high-quality results.
2.2 Dynamic Gaussian Reconstruction and Physics-based Fluid Simulation
The neural radiance expression of objects with dynamics has long been studied for NeRF systems, including deformation capture by canonical and displacement fields  [PSB21, PSH21] and dynamics 3D synthesis  [GSKH21, LNSW21a, LCM22, PCPMMN21, QGX23]. Similarly, for the later proposed 3D Gaussian Splatting, dynamics properties have been imposed on the explicit radiance representation of Gaussians. Leveraging the high fitting ability of Gaussian particles, many studies have explored dynamic scene reconstruction guided by image-based losses from video [RXM24, WYF24, YPTW24, GXC24, NRS22, DWD24].
In physics-based simulation, fluids are often simulated using Eulerian grids and/or Lagrangian particles [Bri15], with different consideration of viscosity  [BT07, DG95]. Methods of solving these equations have advanced with numerical discretization schemes  [FM96, TM94]. For simulation stability, Stam  [Sta99] introduces the concept of semi-Lagrangian advection which brings up the idea of the hybrid field. Hybrid methods  [Har88, BR86, JSS15, FHNJ20, JST16, HFG18] combine advantages of Lagrangian and Eulerian schemes, representing fluid by particles while computing dynamics on grids.
Several studies have augmented static Gaussian points with physical parameters to make them animatable. PhysGaussian [XZQ23] integrats MPM framework to enable reconstructed 3D Gaussians with versatile dynamic behaviors, and GaussianSplashing [FFS25] combines Position-based Dynamics with advanced rendering methods to represent Gaussian particles with fluid dynamics and appearances. PhysMotion [TJL24] proposes a framework for animating 3DGS generated from a single-view image. The sequential rendering results are then composited with the input background through an inversion and diffusion process to obtain videos. Instead of giving physical parameters to reconstructed Gaussian subjectively, several works have developed the learning procedure of 3DGS dynamics from references. Simplicits [MSP24] uses a neural field of learnable weights for reduced-order simulation. These works mainly apply to unsupervised simulation. Given the guidance, PhysDreamer [ZYW24] optimizes material parameters through a differentiable MPM simulation to reproduce dynamic Gaussians that behave similarly to input videos. BAGS [PTZ24] learns the weights of imposed Gaussian Ellipsoid Neural Bounds to animate the reconstructed object according to the input videos.
In our work, we embed generated Gaussians with physical parameters compatible with APIC method, to align the dynamic behavior of simulated Gaussian particles closest to that of the fluid in the video.
2.3 Velocity Extraction from Videos
We limit the discussion to fluid velocity reconstruction from input videos and images, instead of direct velocity data or experiments. To collect pixel information from videos, optical flow could be effectively utilized in neural reconstructions  [LNSW21b, DZY21] for general flowing scenes that vary for image styles and object appearances. Targeting on specific fluid textures can noticeably improve learning of dynamics and re-simulation accuracy  [GDWY22, LCN23]. Generally, multi-view inputs are required for 3D velocity reconstruction and novel view fluid synthesis relies on at least sparse viewpoints of the fluid for trained neural networks to infer the unseen sides  [ODAO15, EHT18, EUT19]. Thus, previous work seldom focuses on reconstructing the entire velocity field from the single input as a generative process. To inform the extracted velocity field with physical grounded features, DVP [DYWZ23] successfully encodes vortex features to learn the specific eddy dynamics of fluid but is restricted to 2D profile. Other works  [FST21, CLZ22, YZG24] extend to 3D space with strict physical constraints and volume rendering, achieving high-fidelity reconstruction while limited to certain fluid types. Recently, FluidNexus [GYZW25] has employed the novel-view video synthesizer to reconstruct the velocity field that closely match input frames. In our work, we use single-view images to resimulate the 3D velocity field available for arbitrary perspectives, leveraging the idea of image-conditioned 3D generation.
3 Method
3.1 Overview
Our goal is to extract and re-simulate fluid physics from single-view videos, which are easily accessible online. To achieve this, we adopt single-view generative 3D methods for geometry reconstruction, with point-based generation being particularly suitable for capturing dynamic properties. This motivates our use of 3D Gaussian Splatting (3DGS) and the integration of current generative 3DGS models. The resulting point representations are compatible with most particle-based fluid simulations. While the open-system nature of fluids makes direct point optimization challenging, our method employs a widely used grid-to-point transformation that combines Eulerian and Lagrangian perspectives, thereby simplifying dynamic optimization with fixed grid coordinates.
Our method takes a single-view video of fluid as input and outputs a 3D fluid physics asset. We design a two-stage pipeline to automatically extract crucial information, like the fluid’s geometry, appearances, physical properties, and motion, from the input video to form the final fluid physics. Stage 1: Geometry and Motion Reconstruction. We employ image-conditioned 3D generative models to generate 3DGS for the input frames and preprocess the generated 3DGSs to improve quality – making them unified, denoised, and dense throughout the volume. With the 3D geometry information, we estimate the volumetric velocity field of the fluid using velocity-free projection from the surface velocity extracted from the frame images (Sec. 3.3). Stage 2: Simulation Parameter Optimization. To estimate the fluid dynamics in the video through physical evolution, we convert the velocity field on points to a grid-based representation to enable the differentiable grid velocity computation and guide the optimization of simulation parameters, including fluid physical properties and boundary conditions (Sec. 3.4).
Figure 1: Our framework consists of five stages: (1) generating 3DGS representation from input frames and preprocessing the 3D Gaussians, (2) estimating 2D screen-space velocities using optical flow with mainstream correction, (3) combining with depth information to obtain 3D velocities while extracting terrain geometry, (4) optimizing fluid properties through differentiable simulation, and (5) post-processing for final rendering.
3.2 Background
3.2.1 Generative 3D Gaussian Splatting
3D Gaussian Splatting is an explicit radiance-based representation of 3D objects, utilizing high-degree features of shape and color for multi-view synthesis. Due to its differentiable volume rendering capabilities, it can be effectively employed in image-to-3D object generation, enabling the alignment of conditioned image textures. This generation procedure, denoted as $G$, can be described as: $G:\{C\}_{w,h}\rightarrow{\{x,\sigma,A,F\}}_{p}$, where $\{C\}_{w,h}$ represents pixels of the conditioned image at position $(w,h)$. $x$, $\sigma$, $A$, and $F$ denote the position, opacity, covariance matrix, and spherical harmonic features of each Gaussian particle $p$, respectively. Ideally, the input image could be recovered through discretized splatting rendering: $\sum_{i\in\mathcal{N}}\alpha_{i}\boldsymbol{SH}(r|_{w,h};F_{i})\prod_{j=1}^{i-1}(1-\alpha_{j})\rightarrow\{C\}_{w,h},$ where $\alpha_{i}$ is the product of $\sigma_{i}$ and the projected 2D Gaussian density of where the kernel intersects with the ray in direction $r|_{w,h}$ from the specific pixel at $(w,h)$, and $\boldsymbol{SH}$ denotes the color calculated with features $F_{i}$. This is equivalent to computing a depth-and-opacity weighted average color of particles along the ray direction. Consequently, individual Gaussian points possess adequate geometric and chromatic information, facilitating their representation of point-based models, such as fluids.
3.2.2 Affine Particle in Cell Method and Dynamic 3D Gaussians
Affine Particle-in-Cell (APIC) [JSS15] is a numerical technique for simulating particles in continuous media, enhancing stability and precision through additional computation of local affine velocity fields. In a time step $n$, the particle velocity is first transferred to the grid together with its affine part:
$(m\vec{v})_{i}^{n}$ $=\sum_{p}w_{ip}m_{p}[\mathbf{v}^{n}_{p}+\mathbf{C}_{p}^{n}(\mathbf{x}_{i}-\mathbf{x}_{p}^{n})]$ (1)
$\vec{v}_{i}^{n}$ $=(m\vec{v})_{i}^{n}\big/\sum_{p}w_{ip}m_{p},$
where $i$ and $p$ denote grid and particle coordinates respectively. $m$ is the mass, $\vec{v}_{i}$ is the grid velocity and $\mathbf{v}_{p}$ is the particle velocity. $\omega$ is the weight function, such as quadratic B-Spline kernel, and $\mathbf{C}$ is the affine matrix associated with each particle. Assuming the simulated media is incompressible, the velocity change $\frac{\partial\vec{v}}{\partial t}$ can be computed on the grid by the following equation:
$\frac{\partial\vec{v}}{\partial t}=$ $-\frac{\nabla p}{\rho}+\nabla\cdot(2\nu\textbf{S}_{ij})+g$ (2)
$\textbf{S}_{ij}=$ $\frac{1}{2}[\nabla\vec{v}+(\nabla\vec{v})^{\mathsf{T}}],$
where $\nabla p$ is the gradient of pressure, $\rho$ is the density, $\nabla\cdot$ is the divergence operator, $g$ is the acceleration of body forces like gravity. $\nu$ denotes kinematic viscosity, modeling the viscous stress by the strain rate $\textbf{S}_{i,j}$ which is the deviatoric part of the velocity gradient. Here the splitting scheme is often employed to decouple this partial differential equation (PDE) for robust and efficient solve [Cho67]. The body force is initially applied explicitly, followed by projecting the velocity to be divergence-free by solving the Poisson Pressure Equation. Subsequently, the viscous stress is computed from the strain rate. We refer the readers to [Bri15] for more details. The updated grid velocity is then transferred back to the particles as states of the next time step $n+1$, and the corresponding affine matrix is simultaneously updated:
$\mathbf{v}_{p}^{n+1}$ $=\sum_{i}w_{ip}\big(\vec{v}_{i}^{n}+\big(\frac{\partial\vec{v}_{i}}{\partial t}\big)^{n}\Delta t\big),$ (3)
$\mathbf{C}_{p}^{n+1}$ $=\frac{4}{\Delta x^{2}}\sum_{i}w_{ip}\mathbf{v}_{i}^{n+1}(\mathbf{x}_{i}-\mathbf{x}_{p}^{n})^{T}~(\text{quadratic kernel}).$
Finally, the convective effect of simulated media is achieved on the particles by $\mathbf{x}_{p}^{n+1}=\mathbf{x}_{p}^{n}+\mathbf{v}_{p}^{n+1}\Delta t$.
3.3 Reconstruct 3D Gaussians with Dynamics
3.3.1 Preprocess with Generative Gaussian Splatting
Since our input only contains single-view information, to obtain the 3D information, we rely on the existing single-image to 3DGS generation methods [ZYG24, XLX24]. However, these methods are trained on datasets of 3D surface representation and the generated 3D Gaussians are typically sparse and concentrated on the object’s surfaces. The generated 3DGS need to be processed so that the Gaussian points could precisely carry every pixel information from the frame.
The directly generated 3DGS exhibits smooth appearance variations across different viewpoints but suffers from coarseness due to input image compression as shown in the inset left. Large elliptical shapes are thereby produced, which is undesirable when relating continuous surface velocities with such sparse 3D Gaussians. To address this, we perform a fast single-view optimization using frames from the input video to enhance the resolution of the reconstructed 3DGS (Fig. 2) While using only one viewpoint significantly reduces computational time, it necessitates a carefully designed loss function to mitigate overfitting due to limited supervision:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{img}+\lambda_{2}\mathcal{L}_{aniso}+\lambda_{3}\mathcal{L}_{vol}+\lambda_{4}\mathcal{L}_{scl}+\lambda_{5}\mathcal{L}_{lumi}.$ (4)
$\mathcal{L}_{img}$ denotes the image loss adopted from original 3D Gaussian Splatting [KKLD23]. $\mathcal{L}_{aniso}$ (anisotropic loss) and $\mathcal{L}_{vol}$ (volume loss ) are implemented the same as Gaussian SPlashing [FFS25]. $\mathcal{L}_{scl}$ represents the geometric mean of Gaussian ellipsoids’ three scale components. This term encourages finer particle details during optimization. $\mathcal{L}_{lumi}$ is the luminance loss, penalizing oversaturated colors in harmonic features. During optimization, we prune Gaussians whose colors closely match the background. This strategy effectively eliminates the misrepresentation of rocks or reefs near the water surface. Meanwhile, this optimization uses a lower threshold than the original 3DGS to encourage densification.
Figure 2: Preprocessed results for generated 3DGS. The left column is the direct generation, and the right column is processed with the single-view optimization.
Subsequently, the optimized 3DGS are denoised and densified. The Gaussian primitives with low opacity or heavily stretched covariance are first pruned. As shown in the inset, we then sample the 3DGS space into grids and insert a Gaussian at the center of the grid $p_{c}$ if $p_{c}$ is inside the fluid by checking if over half of the number of intersections of 3DGS’s outer hull and random rays from $p_{c}$ is odd.
We also need to guarantee the consistency of geometry for the generated 3D Gaussians in continuous frames. Generative methods do not guarantee that the output geometry varies as continuously as input images. Moreover, adjacent frames could be generated with quite different geometries at the backside from the camera, though the foreground is conditioned by similar inputs. This is acceptable for free-moving fluid like smoke, which shows highly dynamic behavior in the video. However, this inconsistency can impact the calculation of physically grounded fluid dynamics for flowing rivers on fixed riverbeds. To address the issue, we perform the union operation on the generated 3D Gaussians from $N$ consecutive frames, forming a batch to be used in the subsequent process of simulation parameter optimization in Sec. 3.4. The frame number $N$ is determined dynamically based on the motion intensity of fluid objects in the video. This is achieved by evaluating similarities between adjacent frames $f$ and $f+1$: $N\propto MSE\big(PSNR({f,f+1})\big)$. The effects of filling and union are seen in Fig. 3. Filling aims to fill in the inner vacancy of the generated fluid body that prevents unreal collapse in the simulation. Union mainly aims to make fluid geometry consistent across frames, especially for the riverbed, which is supposed to be invariant during the flow motion.
Figure 3: The filling operation inserts Gaussians into sparse generated 3DGS in (a) and output a dense 3DGS in (b). Our union strategy merges the generated 3DGS from multiple frames and outputs a higher-fidelity geometry in (c).
3.3.2 Fluid Surface Velocity Estimation
After obtaining the geometry, we can estimate the volumetric velocity field $V_{vol}$ of the object. This starts by first estimating the 3D fluid surface velocity $V_{surf}$, where we first utilize optical flow [Far03] to detect the 2D velocity field in the screen space, denoted as $V_{2D}$, as pixels moving on the screen can be seen as material points being displaced in the Normalized Device Coordinates (NDC). Additionally, we combine the $z$-depth information from the preprocessed 3DGS to get displacements in the third dimension. However, optical flow fails to detect the velocity $V_{2D}$ in locally homogeneous fluid textures, often yielding zero velocity. The detected pixel motion often corresponds to waves or splashing – waves partially reveal the main flow direction, while splashing provides only local information.
To tackle this problem, we propose a mainstream correction strategy and compensate the region where optical flow fails. The process comprises two stages: (1) mainstream-guided neighboring interpolation and (2) physics-constrained velocity correction. Stage (1): In regions where optical flow detection fails, the underlying motion typically exhibits smoothness and reflects bulk fluid behavior. We determine the mainstream direction using geometric constraints (e.g., river bank alignment) and estimate its magnitude from neighboring optical flow, which ought to have comparable kinetic energy.
$\vec{v}_{opz,k}=\sum_{i\in B(k)}w(i)\max(0,\bigg(\vec{n}_{k}\cdot\frac{\vec{v}_{opz,i}}{|\vec{v}_{opz,i}|}\bigg))\cdot\vec{v}_{opz,i}.$ (5)
$\vec{v}_{opz}$ is the 3D velocity in NDC space, calculated from optical flow and depth change. $\vec{n}$ is the mainstream direction, $i$ denotes the detectable place by optical flow, and $k$ is the missing part. $B$ bounds the range of $i$ contributing to specific $k$ and defines a distance-based weight $\omega(i)$. The dot product of mainstream direction $\vec{n}_{k}$ and the normalized $\vec{d}_{i}$ measures the cosine value of their angle.
Unlike elastic bodies, interpolation alone is unsuitable for the amorphous features of liquids. In Stage (2), we derive constraints, based on the divergence-free property of the 3D velocity $V_{surf}$, that the 2D screen velocity $V_{2D}$ must satisfy:
$\nabla_{2D}\cdot\vec{v}_{2D}=-\frac{1}{z}(u\frac{\partial}{\partial u}+v\frac{\partial}{\partial v}+2)v_{z},$ (6)
where $u,v$ are screen coordinates, and $z$ is the depth. The detailed derivation can be found in the Appendix (Section 1.1). This constraint depends exclusively on the out-of-plane velocity $v_{z}$.
In regions where optical flow fails to detect the motion, this constraint cannot uniquely determine the 2D velocity $V_{2D}$ because the optical flow data does not form a closed and well-posed boundary. To resolve this, we employ the projection method with $V_{2D}$ initialized in the aforementioned mainstream direction. We apply this mainstream correction only to areas where optical flow fails, thereby preserving existing turbulence. Based on this interpolation, we then project $V_{2D}$ to satisfy the constraint in Eq.6. Fig. 4 shows that the mainstream-corrected result effectively captures the true motion where the original optical flow detects near-zero velocities, and interpolation alone leads to noisier dynamics.
Finally, the 2D velo city $V_{2D}$ can be inversely mapped to the 3D Gaussian space according to the camera parameters $\vec{v_{surf}}=\big(S\circ(PW)+T\big)^{-1}\circ\vec{v}_{2D}$ , where $\vec{v}_{surf}$ is the 3D velocity of surface particles, $P$ is the view projection matrix, and $W$ is the world-to-camera matrix. $T$ and $S$ denote the transition and scaling same as the preprocess (if any) done on input images before generating the 3DGS. This approach yields physically plausible surface velocities $V_{surf}$ for subprocesses.
Figure 4: The unguided result shows vanished velocity in regions such as the one highlighted in the close-up (left). After applying mainstream-guided neighboring interpolation, we obtain the result shown in the middle. With further physics-constrained velocity correction, we achieve a more meaningful velocity field (right).
3.3.3 3D Volumetric Velocity Estimation
With the surface velocity $V_{surf}$ extracted, we calculate the entire volumetric velocity $V_{vol}$ of preprocessed Gaussian particles to initialize the dynamic behavior of the fluid object. Velocity on the enclosure surface, $V_{encl}=V_{surf}\cup V_{base}$, needs to be fully determined before computing $V_{vol}$. The surface velocity $V_{surf}$ from the previous stage is used, with additional unknowns $V_{base}$ at the riverbed. While the analytical solution requires zero velocity near solid boundaries, discretization allows for damping, meaning that the velocity at the grid cells closest to the boundary does not vanish in the simulation. We employ wall functions from computational fluid dynamics (CFD) to estimate this damping, which depends on grid resolution and fluid type. Although this assumes external laminar flow—limiting fine details—it avoids introducing nonphysical supervision in later optimization. This ensures a gradual change in $V_{base}$ of the flowing layer as it approaches any solid boundary, as shown in Fig. 5. To reveal this effect, we choose the third-order polynomial of the integral approximate solution of laminar external flows:
$\frac{|\vec{v}_{base}|}{V_{\infty}}=\frac{3}{2}\left(\frac{y}{\delta}\right)-\frac{1}{2}\left(\frac{y}{\delta}\right)^{3},$ (7)
where $V_{\infty}$ is the magnitude of corresponding surface velocity, $\delta$ is the boundary layer thickness, which is dynamically decided by the fluid type: for a flowing river it is set to a larger thickness, while for smoke, it is limited to a sub-cell scale so it introduces very small damping. $\vec{v}_{base}$ takes the direction of the mainstream but not the same as its surface counterpart.
Figure 5: Illustration (left) and a real volumetric velocity field (right). The velocity is dampen near the riverbed. On the right figure, the darker the color, the smaller the velocity.
After obtaining the enclosure surface velocity $V_{surf}$, we use a constrained projection method (Eq. 8) to solve for a divergence-free 3D velocity field, with the aforementioned boundary correction providing a closer initialization.
$C_{i}:~(\nabla\cdot\vec{v})_{i}=0,~~~~~~~~~\vec{v}_{j}^{n+1}=\vec{v}_{j}^{n}+\Delta\vec{v}_{j},$ (8)
$\Delta\vec{v}_{j}=\Sigma_{i}\lambda_{i}\nabla_{v_{j}}C_{i},~~~~~~~\lambda_{i}=-\frac{C_{i}(\vec{v}_{1},\dots,\vec{v}_{n})}{\Sigma_{k}|\nabla_{v_{k}}C_{i}|^{2}+\epsilon}.$
For each grid cell $i$, the divergence-free constraint $C_{i}$ depends on its neighboring velocities $\vec{v}_{k}$. During each iteration of the constrained projection solver, the velocity update $\Delta\vec{v}_{j}$ is computed based on the gradients $\nabla_{v_{j}}C_{i}$ of all constraints influenced by $\vec{v}_{j}$. The relaxation parameter $\epsilon$ ensures stable convergence.
3.4 Presumed APIC Simulation and Optimization
3.4.1 Grid-Based Velocity Evolution
We transfer the velocity field from particles to the grid, which offers two main advantages. First, grid-based representation helps us more efficiently track the fluid movement between two frames. Second, by solving Eq. 2 alongside grid-based convection, particle velocity can be transferred to grid using Eq. 1 with $\mathbf{C}$ set to zero, making the dynamics of Gaussians differentiable (see Appendix, Section 2.1 for details).
To formulate the optimization problem , we modify two parts of Eq. 2. First, to solve the Poisson Pressure Equation, instead of using iterative methods, we implement and modify the PeRCNN [RSL21] structure shown in the inset where $P^{(k)}$, $B.C.$ and $\vec{v}$ are the latents of pressure, boundary condition and velocity respectively at $k$-th recurrent layer. PeRCNN replaces the implicit pressure solver during optimization. As demonstrated by PDE-Net [LLMD18], a convolutional kernel with properly designed weights can effectively discretize spatial operations equivalent to those in numerical methods. Different weight configurations can approximate various differential operators (e.g., gradient, divergence, Laplacian), and larger stencil sizes typically enable higher-order discretization schemes. Beyond the convolutional structure, an outer recurrent layer is formed to emulate the iterative process of implicit solvers. This leverages the RCNN structure to facilitate optimization while maintaining effective pressure correction, without requiring an ad-hoc PeRCNN model.
We employ this architecture to solve the Poisson pressure equation in our physical optimization framework, leveraging its differentiable properties and improved stability for convergence. Prior to implementation, we first validate the structure’s numerical accuracy and computational efficacy. This is done by sampling data from our reconstructed 3D velocity field with initialized physical properties, recording the pressure field both before and after solving via an implicit solver. These paired data samples are then used to evaluate the PeRCNN, which is expected to predict pressure results the same as the second-order Laplacian discretization scheme used in the numerical method. As shown in Fig. 6, the validation demonstrates successful convergence; we consequently employ the fitted kernel to solve pressure in subsequent physical optimization, whose efficacy has been demonstrated in the paper.
Figure 6: Optimization network (left) and parameter fitting of its convolutional kernel for pressure solving.
Second, as the original APIC simulation method does not solve the convection part on the grid, we additionally integrate the back-tracing mechanics from stable flow [Sta99] in the forward process, allowing for unconditional stability in resolving grid-based velocity. However, since back-tracing involves grid indexing and round-off operations, which reduce the differentiability of the optimization process, we directly use the analytic gradient of the convection term of the split PDE: $\frac{\partial\vec{v}}{\partial t}=-\vec{v}\nabla\vec{v},~\frac{\partial\vec{v}}{\partial v_{i}^{\prime}}=-dt\frac{\partial\vec{v}}{\partial i},$ where $\vec{v}$ denotes the grid velocity after convection, $v_{i}^{\prime}$ denotes the velocity component of dimension $i$ before convection. For first-order spatial discretization, we employ the upwind scheme to enhance stability.
Fig. 7 demonstrates that our modified PeRCNN [RSL21] architecture for solving the Poisson pressure equation is equivalently accurate compared with traditional numerical methods using iterative solvers or analytical gradients. Also, the comparison between the optimization processes of PeRCNN and Jacobi iterations reveals that PeRCNN achieves faster convergence.
Figure 7: We compared the optimization process whose Poisson pressure equation is solved by PerCNN structure with those solved by (left) Jacobi iteration and (right) analytical gradient. The latter two methods apt for different batch sizes. Both comparisons demonstrate that utilizing the PeRCNN structure leads to faster convergence.
3.4.2 Loss Design and Parameter Optimization
Prepared with the differentiable computation of grid velocity, we can optimize several parameters to provide preprocessed 3D Gaussians with physical properties, including bulk velocity at the inlet $v_{in}$ and outlet $v_{out}$, velocity fluctuations at the inlet $\tilde{v}_{in}$, density $\rho$ and viscosity $\nu$ of the fluid, bouncing $b$ and damping $d$ coefficients at the boundary, gravity $g$, and the time step $dt$ of the simulation. To design a reasonable loss function for the optimization, we combine the $L_{2}$ loss of grid velocities and the dot product of normalized velocities as the final loss, masked by grid types of surface, occupation, or empty.
$\vec{v}_{sim}^{i+n}=\text{DiffSim}^{n}$ $\big(\vec{v}^{i}_{gt};v_{in},\tilde{v}_{in},v_{out},\rho,\nu,(b,d)|_{{}_{B.C.}},g,dt\big)$ (9)
$\min~\text{gridMask}\circ$ $\bigg(\alpha(1-\frac{\vec{v}_{sim}}{|\vec{v}_{sim}|}\cdot\frac{\vec{v}_{gt}}{|\vec{v}_{gt}|})+\beta L_{2}(\vec{v}_{sim},\vec{v}_{gt})\bigg).$
During the experiment, each training sample could require up to 25 continuous steps, with about 100 recurrent sub-steps per simulation step, creating a deep and lengthy computation graph. Gradient vanishing would occur in cases like zero initialization, convection schemes requiring indexing, even longer optimization steps, or large surface masks. To address this, optimization is carefully designed with reasonable initialization (constant in/out velocity, non-zero gravity, unity relaxation, CFL-satisfying time step - applicable to multiple fluids with similar scales) and upwind gradient surrogates in convection schemes. Step lengths are also adjusted based on scene dynamics (longer for steady-state, shorter for highly dynamic fluids), and results improve with a higher surface/volume ratio in prior 3D velocity recovery. Physical parameters must be normalized during training. Each parameter is rescaled through an activation function (either exponential or sigmoid, depending on whether negative values are permissible) and then mapped to its corresponding physical scale. This normalization enables more balanced learning rates across parameters.
3.4.3 Augmentation of Optimized Result
In the post-processing of the generated 3D Gaussians and optimized simulating parameters, we can re-simulate the Gaussians using the APIC method, combining Lagrangian and Eulerian spaces. When integrating the APIC with Gaussian points, the simulation properties and Gaussian properties are in two-way coupling. On one hand, Gaussian points determine the initial position and possibly the velocity of the simulating particles. On the other hand, the calculated affine matrix records the local deformation of the material and updates the Gaussian covariance matrix:
$F_{p}^{n+1}$ $=(I+dtC_{p}^{n+1})F_{p}^{n},$ (10)
$A_{p}^{n}$ $=F_{p}^{n}A_{p}^{0}F_{p}^{T}.$
Here we approximate the affine matrix $C$ as the velocity gradient and update the deformation gradient $F$ each time step, which then deforms 3D Gaussians and is reflected by the change of its covariance $A$. An example of the re-simulation of our final output geometry and physics is shown in Fig. 8.
Figure 8: The middle figure shows the re-simulation of our final output geometry and physics based on the input video on the left, which is then rendered in the figure on the right.
3.5 Result Rendering
The output 3D Gaussians dynamic assets can be rendered mainly in two ways: retaining the originally generated spherical harmonic features or reconstructing meshes for high-resolution rendering and interaction. Without assigning uniform appearances to fluid, 3D Gaussian rasterization relies on the unknown spherical harmonic features of inlet particles. Thus, a natural choice is to sample features from particles at the initial position. We note that the generative method often produces Gaussians at low resolution, which may limit the application of harmonic features, depending on the fluid type. Through the experiment, rendering directly by splatting is best suited for viscous fluids or fluids with uniform appearance. The former type of fluid flows more steadily like a deformable soft body, and the latter type is free of appearance inconsistency in fluid motion.
Unlike a closed system, our fluid assets include inlet particles, which complicates the prediction of inflow textures. While the appearance of the fluid typically aligns with uniform flow, it may lack consistency with the reconstructed appearance. To enable interaction between the generated Gaussians and other assets, users might need to re-render the fluid using the same illumination pipeline as the rest of the scene. For compatibility with splatting rendering, we could adapt the existing 3D Gaussian relighting frameworks [JTL24] and implement the Gaussian fluid with transparent light absorption  [FFS25], ensuring rendering efficiency with 3DGS. Moreover, with current GPU-accelerated parallelization, the simulation process achieves real-time speeds, enabling both rasterization and simulation of Gaussians on the fly.
Figure 9: Examples of our output. Left column: frames from the input videos. Middle column: frames from the rendered output. Right column: side views of the output fluid, captured from camera angles indicated by the yellow arrows in the middle images. The last two rows correspond to viscous fluids and they are directly rendered by splatting.
3.6 Editable and interactive features
Unlike 4D reconstruction methods based on canonical and displacement fields, recovering fluid dynamics using a physics-based approach offers enhanced editability and interaction with the generated geometry. An optimized set of physical parameters provides an explicit representation of fluid motion, making it more interpretable for artists and users without specialized expertise. In contrast, directly optimized displacement fields typically possess a high degree of freedom, which complicates the interpretation of their influence on fluid behavior. In our method, any modifications to the optimized parameters or external force interactions are propagated through fluid re-simulation, ensuring all changes remain physically consistent. This capability is particularly advantageous when scene reconstruction is not the sole objective (e.g., when extracting 3D assets from real-world videos and integrating them into digital scenery). Moreover, our approach not only recovers the fluid body but also generates the surrounding and underlying terrain, providing enriched geometric information for diverse post-processing applications.
4 Results
We implement the framework of our method mainly using Python on a workstation PC equipped with a 24-core Intel(R) Xeon(R) Platinum 8255C CPU and an NVIDIA Tesla V100 GPU. For the 3DGS generation part, we port the open-source implementation of the existing methods (we use TriplaneGaussian [ZYG24] by default), and output intermediate states. For re-simulating the flowing fluid, we implemented a GPU-parallelized APIC method with Taichi [HLA19]. The rendering of obtained 3D Gaussians is compatible with a general 3DGS visualizer. We also implement a relighting interface using GaussianShader [JTL23] to achieve higher resolution and incorporate transparent fluid rendering within the original 3DGS CUDA rasterizer.
4.1 Velocity Optimization and Reconstruction
Figure 10 compares four types of $V_{vol}$: (1) velocity estimated from the input video and used as optimization guidance, (2) a simulation with random physical parameters, and two optimized simulations employing either (3) partial or (4) complete loss terms as defined in Eq. 9. This is also used as an ablation study to verify both the necessity and efficacy of our optimization approach. Specifically, we examine the extent to which the optimization changes the parameters and how these changes influence the simulation outcomes, particularly in terms of the resulting 3D velocity fields. Besides, as shown in Fig. 10, the reconstructed dynamics can even correct artifacts present in the video-estimated velocity fields through physical simulation.
Figure 10: Visualization of (a) volumetric velocity obtained from the video, (b) simulation result of randomly initialized physical parameters, (c) simulation result of optimized parameters with only MSE loss, and (d) simulation result of optimized parameters with full loss terms.
Table 1 compares the scale variation of several parameters and the the simulation results before and after the optimization process. Starting from arbitrary initializations at plausible scales, the optimized parameters consistently converge to stable values. This convergence is expected, as these values represent the optimum for the given reconstructed terrain and estimated 3D velocity. In contrast, manual estimation of parameter scales often fails to produce simulation results that match the velocity field extracted from the video. As shown in Fig. 10 (b), simulations using the same geometry but without proper optimization exhibit significant fluid overflow beyond the terrain. Furthermore, even initialized with scales comparable to the final optimized values, the subtle discrepancies in the magnitude can lead to pronounced differences in the simulation over time, ultimately making the dynamic reconstruction fail.
[…]
Table 1: Optimization effects on typical physical parameters and simulated results for the reconstruction scale as the case in Fig. 10.
Two synthetic videos generated from known fluid simulation data are used to compare the reconstructed 2D velocity fields against the ground truth. Since our method inputs planar velocity (discretized by frame rate and scaled to screen dimensions), both the ground truth and reconstructed 3D velocities are projected into 2D NDC space for visual comparison. Figure 11 presents the result. The reconstructed in-plane velocity captures mesoscale dynamics but retains noise on fluid surfaces. By optimizing physics-based simulation and aligning velocity directions, we achieve smoother results while enhancing physical details—such as boundary collisions in the river flow and strain development in the honey’s inflow region.
Figure 11: Comparison of planar velocity for fluid as river (top) and honey (bottom). Column (a) is the ground truth, column (b) is the estimated planar velocity in early stage of the framework as described Fig.1, and column (c) is the result by our method.
4.2 User Study
We conducted a user study to perform a qualitative evaluation, comparing the results of our method with those generated by TriplaneGaussian across all frames of the input videos (videos of the results are included in the supplementary material). The study included the 8 examples in Fig. 3D Dynamic Fluid Assets from Single-View Videos with Generative Gaussian Splatting and Fig. 9. Users were asked to evaluate the outputs along three dimensions: (1) similarity to the input video, (2) realism of the output, and (3) aesthetic quality. For each dimension, participants rated the results on a scale from 1 to 5, with higher scores indicating better performance. We invited 15 participants, all of whom are either full-time 3D artists or graduate students specializing in 3D design. The scores from the user study are visualized in Fig. 12, where our method achieves noticeably higher scores overall.
Figure 12: Plot of the scores from the user study. Left: scores for TriplaneGaussian. Right: scores for our results. Higher scores, indicated by the greener regions, represent better performance.
4.3 Evaluation on Different Fluid Types
We apply our method to two different types of liquid fluids – inviscid and viscous liquid – to show the generalization of our framework in Fig. 9. The effects of gravity and boundary conditions are more pronounced in inviscid fluids with well-defined shapes (Fig. 9 (a) to (e)). For viscous fluids (Fig. 9 (f) to (g)), the relative motion between different parts is constrained by viscosity and further dampened by the boundary. By optimizing additional simulation parameters (e.g., viscosity), the resimulated fluid can mimic a wide range of dynamics observed in the video. Meanwhile, the rocks on the water surface (Fig.3D Dynamic Fluid Assets from Single-View Videos with Generative Gaussian Splatting, Fig.9 (a) and (e)) can also be successfully detected, benefiting from the pre-processing of the generated 3DGS.
4.4 Performance with Different 3DGS Generation Methods
As a generic framework based on generative 3DGS, our method is not dependent on a specific method of Gaussian generation. We test our method with different 3DGS generation methods: generative methods, TriplaneGaussian (Fig. 9 (a), (c) to (g)) and TRELLIS [XLX24] (Fig. 9 (b)). These methods differ in terms of image fidelity, resolution, and creativity, thus catering to varying user needs in real-world applications. TriplaneGaussian benefits from triplane encoding of image textures, which aligns well with the input image, though this comes at the cost of lower resolution. TRELLIS shows more divergence from the input images, even at high Classifier Free Guidance (CFG) scales, but produces Gaussians at a higher resolution compared to TriplaneGaussian.
4.5 User Editing
For users looking to edit the generated fluid asset, the framework provides easy modification of simulation parameters to achieve different fluid effects, change the fluid’s material, adjust the downstream flow, or introduce solid-liquid interactions as shown in Fig. 3D Dynamic Fluid Assets from Single-View Videos with Generative Gaussian Splatting and Fig. 13. Our output asset also includes a generated terrain, which makes the fluid easily concatenated to and produce interaction with existing 3D scenery. In this way, fluid behavior within scenic videos can be faithfully reproduced in the digital world.
Figure 13: Examples of our output after user editing. (a): Several interacting objects are added, extending the downstream length of the fluid. (b): The appearance of the fluid is modified, and the gravity direction is changed. (c): Fluid texture can be flexible, generated terrain is twisted, and reconstructed dynamics is interactive.
4.6 Runtime
We analyze the running time consumed by our method. The results indicate that in most cases, simulation parameter optimization dominates the main runtime as the proportion shown in Fig. 14. We note that in the optimization stage, the runtime is influenced by two key factors: the length of the input video and the grid density.
Figure 14: Runtime (in seconds, on a logarithmic scale) of different components of the algorithm for the example shown in Fig. 3D Dynamic Fluid Assets from Single-View Videos with Generative Gaussian Splatting.
5 Conclusions
In this paper, we introduced a novel two-stage pipeline for reconstructing consistent and accurate fluid dynamics from a single-view video, starting with geometry generation and motion reconstruction and then optimizing simulation parameters. Beyond liquid fluids, our method can potentially be extended to gaseous fluids, with necessary enhancements to be explored in future work. Our experiments show that output results of viscous fluids can be directly rendered using the reconstructed 3DGS model. In contrast, for inviscid fluids with more amorphous features, rendering by relighting yields surface textures more consistent with the flow motion. This approach has opened up new possibilities in the field of fluid asset creation. Found by the experiment, our approach currently needs certain adaptions for real use. The quality of the output dynamic 3D fluids would depend on the grid resolution used for optimization and simulation, for input video with flows not that much small-scale dynamics, our appraoch produces satisfcaoty reconsrtuction with afforadble computational cost. but if those videos has complex and detailed splashing, our appraoch needs quite high-res grid-scale and smaller time-step to optimze and simualte, put real-applciation not plausible.
Our experimental results indicate that the practical application of our approach might require certain adaptations. The quality of the reconstructed dynamic 3D fluid dependents on the grid resolution used for optimization and simulation. For input videos dominated by large-scale flow with minimal small-scale dynamics, our method produces satisfactory reconstructions at an affordable computational cost. However, for videos featuring complex, detailed phenomena such as splashing, a higher grid resolution and a smaller simulation time-step are necessary. These requirements can make the computational cost prohibitively high for real-time or interactive applications. Meanwhile, it currently cannot handle long videos with high-torrent fluids due to increased complexity and convergence issues in simulation parameter optimization. Color changes not related to velocity (e.g., in lava due to temperature) may also affect the accuracy of our optical-flow-based velocity estimation before extra parameters are added to the optimization. Future work will aim to address these limitations for more accurate reconstructions.